OpenStudy (anonymous):

The length of a rectangle is 5 inches more than its width, x. The area of a rectangle can be represented by the equation x^2 + 5x = 300. What are the measures of the width and the length?

OpenStudy (anonymous):

Identify what you know and what you do not know: \[Area = length \times width\] \[Length = 5 + x\]\[Width = x\] \[Area = x^2+5x = 300\]

OpenStudy (anonymous):

Solve your problem for x and substitute it back into your length and width.